Activity at the volcano has continued at a low level, with a slight change in the nature of the seismic activity. The visibility was good for most of the day, and a dome survey was carried out. There have been further rockfalls at the Galway's Wall.
The level of seismic activity has been low. A total of 34 volcano-tectonic and 40 hybrid earthquakes were recorded. Some of these occurred in a swarm between 4:22 pm and 6:15 pm. During today, there was a change in the style of activity, with earthquakes occurring throughout the day. Last night's swarm was followed by a short episode of tremor from 6:10 pm to 7 pm. A moderate pyroclastic flow occurred at 6:15 pm, and travelled as far as the Tar River Estate House, resulting in an ash cloud that was blown over Plymouth. Otherwise the level of rockfall and pyroclastic flow activity was low, with only 8 signals recorded.
Visibility was good during the day, and many small rockfalls were observed from Whites. Some of the steep pinnacles on the eastern side of the dome have collapsed since yesterday. A helicopter survey of the dome was carried out, which will help to determine the current growth rate of the dome. The results are still being processed.
A low-amplitude rockfall signal was recorded at 1:20 pm, and this was related to a rockfall over the Galway's Wall. The material from this rockfall did not get as far down the White River valley as previous rockfalls. There has been further erosion of the dome above the top of the wall.
Despite the low level of activity, more pyroclastic flows are expected from the steep eastern face of the dome. Any collapses are likely to be gradual, and produce small or moderate pyroclastic flows. However, a larger and more prolonged collapse can not be ruled out. Residents are reminded to remain alert and responsible at all times, carry ash masks and spend the minimum possible time in the evacuated zone. Zone E, which includes Cork Hill and the airport, remains safe at this time. The Tar River Valley and the upper Galway's area are very dangerous.
Mr Lloyd Lynch, Mr Mark Davies and Mr John Power left the MVO team today. Mr Richie Robertson has resumed his duties as Chief Scientist. Mr Brian Baptie from the British Geological Survey arrived this afternoon for a short visit.